Zayne:

Does any1 understand how to solve points on graphs? Coz I'm confused on what to do

XxXNessalulbaddieXxX:

i gotchu o:

XxXNessalulbaddieXxX:

screen shots?

Zayne:

since im a noob i cant But I can tell you the points. D, E,F

XxXNessalulbaddieXxX:

@zayne wrote:
since im a noob i cant But I can tell you the points. D, E,F
damnn uhhh o-o idkk if i can realy help without an ss

SmokeyBrown:

If you find that you're unable to post a screenshot using the Attach File button, then I would like to request more details about the question. You said there are 3 points, D, E, and F. What are you meant to do with these 3 points; what is the question asking? Also, if the question tells you the coordinates of the points, that would also be useful information to have. All the best to you :)

SmokeyBrown:

@zayne wrote:
it just says Find the perimeter of the triangle in units. Round to two decimal places as necessary.
I see. In that case, you would want to find the lengths of the lines which connect each of the points. That is, find the length of DE, EF, and FD To do that, you may use the Pythagorean theorem, based on the coordinates of each point 3 minutes ago And once you have the lengths of each side, you may add them together to get the total perimeter of the triangle

Zayne:

So how do I put it in the theorium?

SmokeyBrown:

Before we use the theorem, we would need to find the horizontal and vertical displacement between each of the points to get our "a" and "b" values. And to do that, we would need the coordinates of the points. But it sounded like the question did not give you exact coordinates...? In which case, it would not be possible to get an exact answer to the problem. Are you positive that you are not able to take a screenshot and attach it here? Perhaps you can use a Snipping tool or even take a picture of your screen to let us see what the problem looks like?

Zayne:

i am 100% sure that I cannot take a screenshot until i reach 10 smartpoints

SmokeyBrown:

@zayne wrote:
i am 100% sure that I cannot take a screenshot until i reach 10 smartpoints
I had no idea that was a restriction. I guess I can understand why that rule is there, but it's still kind of annoying for us right now. Well, maybe you can describe the points as best as you can, and I can try drawing it out, and then we can go from there?

Zayne:

okay. the points are in a triangle so the graph lines on the right side of the origin goes 5 skips 2 spaces go up 7 and that's where Point D is.

Zayne:

Now on the bottom side, there are 2 Points F, and E Point F is negative 10. connecting to point D at 10.5 Point E is Below -10 on the vertical Y Axis

Zayne:

hopefully that helps

SmokeyBrown:

 Something like this? Or, perhaps I misunderstood your explanation?

Zayne:

that actually looks close to the picture on my assignment tbh

SmokeyBrown:

In that case, I'll assume that the picture matches and continue with the explanation. Even if we're wrong about the coordinates, it should give you an idea of how to solve the problem.  As you can see, we can imagine 3 right triangles such that the 3 sides of our original triangle make up the hypotenuses of the imagined right triangles (drawn in black, blue and green) Based on the side lengths of these drawn-in right triangles, you can use the Pythagorean theorem as discussed above to find the lengths of the hypotenuses (the diagonal lengths), which will tell you the lengths of the original triangle. Once you have the lengths of the original triangle, you can add them together to get the perimeter of the original triangle

SmokeyBrown:

So, for example, if we focus on the black triangle first, we can use the Pythagorean theorem like this: We have two shorter edges equal to 10 and 10; we can use them to find the length of the longer edge 10^2 + 10^2 = c^2 100 + 100 = c^2 200 = c^2 \[c = \sqrt{200}\] Which is about 14.14. So, our first triangle length is 14.14 Would you like to try with the blue and green triangles to find our remaining two triangle lengths?

Zayne:

yeah

SmokeyBrown:

Alright cool. Please feel free to post your calculations and results, and I'll be happy to check them for you

Zayne:

okay thank you